OpenStudy (anonymous):

Find the volume of the pyramid. h=8.5ft 10ft 10ft. A) 425 ft^3 B) 85 ft3^3 C) 850 ft^3 D) 283.33 ft^3 How do I do this problem?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

looks up the formula for the volume of a pyramid, then plug in the numbers

OpenStudy (anonymous):

Hey I was about to tag you lol.

OpenStudy (anonymous):

V=lwh 3

OpenStudy (misty1212):

in google you can just plug in the numbers directly

OpenStudy (misty1212):

put in 10,10,8.5

OpenStudy (anonymous):

K I will try it. Then I will let you know what answer I got.

OpenStudy (misty1212):

\[\frac{10\times 10\times 8.5}{3}\]

OpenStudy (anonymous):

Can i say the answer?

OpenStudy (anonymous):

its 283.33

OpenStudy (anonymous):

Thank you both, so much I got 283.33 ft^3

OpenStudy (anonymous):

No problem